How many atoms are in 56.7 grams Mg3P2?


1
Expert's answer
2020-12-16T05:53:35-0500

"N=\frac{m}{M} \cdot N_A"

where "N" - number of atoms; "m" - mass, g; "M" - molar mass, g/mol; "N_A" - Avogadro's number, "N_A \approx 6.022 \cdot 10^{23} \; mol^{-1}".

"M(Mg_3P_2)=3 \cdot 24.305+2 \cdot 30.974=134.863 \; g/mol"

"N=\frac{56.7}{134.863} \cdot 6.022 \cdot 10^{23} \approx 2.53 \cdot 10^{23}"